OpenStudy (anonymous):

perform the indicated operation and if possible 2/3-1/15 and 9/10 * 6/5

jhonyy9 (jhonyy9):

2/3 -1/15 =(10-1)/15 = 9/15 = 3/5 9/10 * 6/5 =(9*6)/(10*5) = 54/50